OpenStudy (bananas1234):

. Evaluate 3 + √x + 7 - √3x when x = 9.

OpenStudy (bananas1234):

the √ covers x + 7

OpenStudy (academicgurusinc):

Let's do this again!

OpenStudy (bananas1234):

and 3x

OpenStudy (bananas1234):

ok!

OpenStudy (academicgurusinc):

So let me simplify it for you. 3+Sqrt(9+7)-Sqrt(8x9) .... What do you get?

OpenStudy (bananas1234):

3 + 4.123 - 5.196

OpenStudy (academicgurusinc):

Don't get to the decimals! Just tell me in the radical format.

OpenStudy (academicgurusinc):

so we'll have 3+sqrt(16)-Sqrt(72) .... Right?

OpenStudy (academicgurusinc):

Sorry sorry should be 3+sqrt(16)-Sqrt(27)

OpenStudy (bananas1234):

ok.. 3 + √16 - √27

OpenStudy (academicgurusinc):

Yup! Now Simplify!

OpenStudy (bananas1234):

3 + 4.123 - 5.196

OpenStudy (academicgurusinc):

Well... Square root of 16 is just 4! So 3+4-Sqrt(27) ... From there we'll get 7+sqrt(27)

OpenStudy (bananas1234):

ok

OpenStudy (bananas1234):

so 7 + 2?

OpenStudy (academicgurusinc):

We can even go more... Now we can simplify \[\sqrt{27}\] so we'll have \[\sqrt{9*3}\] then we'll get \[7+3\sqrt{3}\] as our final answer :)

OpenStudy (bananas1234):

wouldnt it be 7 - 3√3?

OpenStudy (academicgurusinc):

Actually you right... it's - ... didn't see it!

OpenStudy (bananas1234):

ok np!

OpenStudy (academicgurusinc):

so 7-\[3\sqrt{3}\]

OpenStudy (bananas1234):

OK now i need the same question in decimal form.

OpenStudy (academicgurusinc):

Alright great... So let's go back to our original problem!

OpenStudy (academicgurusinc):

so we got \[7-\sqrt{2}\]

OpenStudy (academicgurusinc):

...... Sqrt(27) i meant

OpenStudy (bananas1234):

ok

OpenStudy (triciaal):